OpenStudy (wade123):

I need someone to double check me please (: Find the sum of the arithmetic sequence. 3, 5, 7, 9, ..., 21

OpenStudy (wade123):

39?

OpenStudy (wade123):

@hartnn

OpenStudy (wade123):

@ganeshie8

ganeshie8 (ganeshie8):

doesn't look correcct

ganeshie8 (ganeshie8):

start by finding out the number of terms in the sequence

OpenStudy (wade123):

11

ganeshie8 (ganeshie8):

try again

OpenStudy (wade123):

10*

OpenStudy (wade123):

sorry miscounted

ganeshie8 (ganeshie8):

last term = 21 : a + (n-1)d = 21 3 + (n-1)*2 = 21 (n-1)*2 = 18 n-1 = 9 n = 10 So there are 10 terms in the sequence

ganeshie8 (ganeshie8):

n = 10 a = 3 L = 21 use the formula : Sum of arithmetic series whose first term = a, last term = L is : \[\large \dfrac{n}{2}(a+L)\]

OpenStudy (wade123):

120 (: @ganeshie8